OpenStudy (anonymous):

solve using subtitution y=-x^2-2x+8 y=x^2-8x-12

OpenStudy (anonymous):

set them equal and solve

OpenStudy (anonymous):

\[-x^2-2x+8 =x^2-8x-12\]

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

then i got them on one side and got -2x^2-6x+20

OpenStudy (anonymous):

hello?

OpenStudy (saifoo.khan):

hen factor and solve.

OpenStudy (anonymous):

thats where i get stuck

OpenStudy (saifoo.khan):

Then*

OpenStudy (anonymous):

are they -10 and 4

OpenStudy (saifoo.khan):

Do you know how to factor ?

OpenStudy (anonymous):

because it has sum of -6 right

OpenStudy (saifoo.khan):

Yes ,you're right

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

i plugged it in and got -2x^2+10x+4x+20

OpenStudy (anonymous):

is this right

OpenStudy (saifoo.khan):

-2x^2 -10x + 4x + 20

OpenStudy (anonymous):

then what?

OpenStudy (saifoo.khan):

Then make pairs. 2 left terms and 2 right terms

OpenStudy (anonymous):

\[-x^2-2x+8 =x^2-8x-12\] don't start with a negative leading coefficient, that makes your life harder \[2x^2-6x-20=0\] then divide by 2 \[x^2-3x-10=0\] factor\[ (x-5)(x+2)=0\]

OpenStudy (anonymous):

ok im kind of lost

OpenStudy (saifoo.khan):

A GPS can help you to get back on track. ;)

OpenStudy (anonymous):

factoring: \[x^2-3x-10\]1) find multiples of -10 who's sum in -3. Which is -5 and 2, then rewrite equation. \[x^2+2x-5x-10\]2) Do some grouping, and simpler factoring\[(x^2+2x)+(-5x-10)=x(x+2)-5(x+2)\]3) Notice anything in common..? We can now rewrite in the simplest factored form\[(x-5)(x+2)\]